The length of a rectangle is 5 inches less than three times its widths. The perimeter is 86 inches. Find the length and width
Answer by algebrahouse.com(1079)

"The length of a rectangle is 5 inches less than three times its widths. The perimeter is 86 inches. Find the length and width."

x = width
3x - 5 = length {length is 5 less than 3 times width}

2x + 2(3x - 5) = 86 {perimeter = (2 x width) + (2 x length)}
2x + 6x - 10 = 86 {used distributive property}
8x - 10 = 86 {combined like terms}
8x = 96 {added 10 to both sides}
x = 12 {divided both sides by 8}
3x - 5 = 31 {substituted 12, in for x, into 3x - 5}

width = 12 in.
length = 31 in.
www.algebrahouse.com